OpenStudy (anonymous):

Find the sum series 4 ∑ k-1 = (k+4)^2

OpenStudy (anonymous):

\[\sum_{k=1}^4(k+4)^2\]?

OpenStudy (anonymous):

yes thats how it looks i need help

geerky42 (geerky42):

\[\Large \sum_{k=1}^{4}(k+4)^2 = (1+4)^2 + (2+4)^2 + (3+4)^2 + (4+4)^2\]

OpenStudy (anonymous):

the problem is cut off at the end

geerky42 (geerky42):

= (1+4)² + (2+4)² + (3+4)² + (4+4)²

OpenStudy (anonymous):

i got 125

geerky42 (geerky42):

Well, I got 174. (1+4)² + (2+4)² + (3+4)² + (4+4)² = 5² + 6² + 7² + 8² = 174

OpenStudy (anonymous):

so do i type the term of the series